# Engineering interaction potentials for stabilizing quantum quasicrystal phases

### v1

@@ III Dodecagonal quasicrystal
Moreover, we investigate how the superfluid properties of the $12$-QC phase are impacted by the properties of the pair interaction potential. We estimate the superfluid fraction by means of the Legget’s criterion and consider the scenario posed by Fig. 1(a). The superfluid fraction is calculated across a vertical line of the phase diagram of Fig. 1(d), specifically we consider $\lambda^{2}U\rho=0.50$ and use $\hat{v}(\tilde{k}_{2})$ as the running parameter. The results are shown in Fig. 2(d), where the blue points corresponds to the values of the superfluid fraction obtained from the density profile calculated using the mean field spectral variational method while the orange diamonds corresponds to the profile determined via the GPE solution. Within the HS1 phase, for $\hat{v}(\tilde{k}_{2})\gtrsim-1.5$, we observe a superfluid fraction essentially constant. This is explained by the fact that Fourier modes with $\lvert\bm{k}\rvert=\tilde{k}_{2}$ are not present in the HS1 state when the characteristic wave vector of the pattern is of the order of $\tilde{k}_{1}$. Moreover, the fact that superfluidity is significant along this phase allows us to conclude that it is actually a supersolid phase. On the other hand, in the region in which the model displays a $12$-QC phase, we can notice how the increase of the depth of the second minimum decreases the superfluid fraction. This occurs because decreasing the energy cost of the second minimum promotes higher Fourier amplitudes which implies in a more localized quasicrystalline density pattern, hindering superfluidity in the system. Nonetheless, despite the tendency of the superfluidity, within the quasicrystalline phase its value is still significant, allowing us to confirm that this phase is actually a superfluid dodecagonal phase.
@@ IV Decagonal quasicrystal
Now we proceed to study the stabilization of the decagonal quasicrystal ($10$-QC) phase. We explore systematically pair potentials with multiple length scales in the form of several minima in $\hat{v}\left(k\right)$, positioned to enhance the stability of this structure. Naturally, the first model to be considered contains two minima, the first one at $k=\tilde{k}_{1}=1$ and the second one at $k=\tilde{k}_{2}$ (see Table 2). The corresponding pair interaction potential is displayed in Fig. 3(a), while the associated ground-state phase diagram is shown in Fig. 3(d). In this case, a phase transition is observed at moderate and high intensity of the pair potential $\lambda^{2}U\rho$ as the vertical position of the high momentum minima is varied. For low enough values of $\hat{v}(\tilde{k}_{2})$, the system display a transition between two hexagonal solid phases in which the characteristic wave vector of the patterns changes from $\tilde{k}_{1}$ (HS1) to $\tilde{k}_{2}$ (HS2). Interestingly, despite the fact that this model produce the excitation of the two main wave vectors of the decagonal quasicrystal, this pair potential configuration is still favorable to the hexagonal structure even when the high momentum minimum dominates. This indicates the necessity of introducing additional minima in $\hat{v}(k)$ to achieve the stabilization of the decagonal quasicrystal pattern.
@@
To proceed we consider the case in which a third minimum in $k=\tilde{k}_{3}$ (see Table 2) is added to a two minima degenerate potential with $\hat{v}(\tilde{k}_{1})=\hat{v}(\tilde{k}_{2})=-1$. The resulting model for $\hat{v}(k)$ can be observed in Fig. 3(b) and the corresponding ground-state phase diagram using as running parameters the $\hat{v}(\tilde{k}_{3})$ and $\lambda^{2}U\rho$ is shown in Fig. 3(e). As can be noticed, the introduction of a third characteristic length scale allows the stabilization of a 10-QC phase for a certain range of $\hat{v}(\tilde{k}_{3})$. Interestingly, if $\hat{v}(\tilde{k}_{3})$ is deep enough the system reenters into a hexagonal phase (HS3) with characteristic wave vector $\tilde{k}_{3}$. Moreover, the quasicrystal phase exists up to the low intensity potential regime, making the system to present a direct transition from the homogeneous to the quasicrystal phase. However, this behavior is only present in a limited region of $\hat{v}(\tilde{k}_{3})$, raising naturally the question about what can be done to enhance further the stability or extension of this phase within the phase diagram.
@@
Figure 4: (a) Real space mapping for the decagonal structure in the phase diagram of Fig. 3(f) for parameter values $\hat{v}(\tilde{k}_{4})=-1.50$ and $\lambda^{2}U\rho=0.50$ and its corresponding (b) diffraction pattern. The red circles indicates the two minima and the size of the Fourier modes in blue disks is scaled by a linear function, where the zero harmonic mode has been omitted. (c) Most relevant energy contributions to the total energy of the quasicrystalline structures as a contraction in the $\tilde{k}_{1}=1$ layer for the decagonal vectors. (d) Superfluid fraction for a fixed density value $\lambda^{2}U\rho=0.50$ in the decagonal quasicrystal for the ground state depicted in Figure 3(f). The blue points represents the superfluid fraction obtained via mean field spectral variational method while the orange diamonds are determined via GPE simulations.
@@
In order to answer this question, it is natural to consider the introduction of an additional length scale in the form of a minimum in $\hat{v}\left(k\right)$ at the remaining characteristic wave vector of the decagonal quasicrystalline structure. In this way, we add a minimum at $k=\tilde{k}_{4}$ (see Table 2 ) to a three minima potential that does not display a quasicrystal phase at any intensity ($\lambda^{2}U\rho$) of the pair potential. In terms of minima position, this model satisfy $\hat{v}(\tilde{k}_{1})=\hat{v}(\tilde{k}_{2})=-1$, $\hat{v}(\tilde{k}_{3})=0$ and $\hat{v}(\tilde{k}_{4})$ is taken as running parameter for the construction of the corresponding phase diagram. The typical form of this kind of pair potential is displayed in Fig. 3(c), while in Fig. 3(f) we present the corresponding ground-state phase diagram obtained. As can be observed, when the depth of the fourth minimum at $\tilde{k}_{4}$ is increased the system transitions from a HS1 to a $10$-QC phase. This means that only the addition of this extra minimum promotes the stabilization of the $10$-QC phase, but the extension of the quasicrystal phase is greatly enhanced.
@@
To deepen our understanding about the $10$-QC phase, we now focus on its structural properties. In Fig. 4(a) and (b), we present a typical density pattern configuration and its corresponding Fourier transform for a pair potential of the type presented in Fig. 3(c) ($\hat{v}(\tilde{k}_{4})=-1.50$ and $\lambda^{2}U\rho=0.50$). The radius of each circle in the Fourier transform plot is proportional to the corresponding Fourier amplitude. As we can see, the main modes excited in the density pattern corresponds to those promoted by the minima structure of the pair potential as in the case of the $12$-QC phase. An analysis of the potential energy contractions leading to dominant contributions, analogous to the one performed in the case of the dodecagonal structure, reveals that this structure does not have exclusive amplitude contractions which are specially impactful as in the previous case considered. In this case, contractions allowed for all quasicrystal structures, such as the one depicted in Fig. 4(c) for the decagonal pattern, are the leading contribution in the potential energy expansion.
@@
The behavior of the superfluid fraction for the model in Fig. 3(c), keeping $\lambda^{2}U\rho=0.50$ and using $\hat{v}(\tilde{k}_{4})$ as our running parameter, is shown in Fig. 4(d). We can observe that within the region corresponding to the HS1 phase, modifying $\hat{v}(\tilde{k}_{4})$ does not affect the superfluid fraction. As previously observed, since $\tilde{k}_{4}$ does not match any resonant momentum of the hexagonal crystal solution with characteristic momentum $\tilde{k}_{1}$, the hexagonal ground state is not affected by variations of $\hat{v}(\tilde{k}_{4})$ and hence neither the superfluid fraction. On the other hand, when the system transitions to the decagonal phase, we observe that making deeper the minimum at $\tilde{k}_{4}$ results in a decrease of the superfluid fraction. This behavior is a consequence of increasing the stability of the quasicrystalline phase, which results in an increase of the quasicrystal Fourier amplitudes. This in turn reflects on a more localized density profiles which hinder superfluidity. Moreover, is interesting to notice that close to the quasicrystal phase boundary, the superfluid fraction takes significant values, indicating the existence of a superfluid decagonal phase.
@@
Figure 5: (a)-(c) Pair interaction potential minima structure in Fourier space for the corresponding phase diagrams in (d)-(f), where the first length scale is positioned at $k=\tilde{k}_{1}=1$ and the secondary resonant scales are determined in Table 2. Phase diagram for the minima structure associated with the octagonal quasicrystal where (d) $\bar{k}=\tilde{k}_{2}$ (e) $\bar{k}=\tilde{k}_{3}$ and (f) $\bar{k}=\tilde{k}_{4}$ indicates the position of the minimum in reciprocal space for which we modify the depth as a parameter to determine the ground state of the system. The notation employed to label the phases is analogous to the one established in Fig.1.
@@ V Octagonal quasicrystal
Finally, we focus on the study of the stability of the octagonal quasicrystal ($8$-QC) phase. As in the previous cases, we initially consider a two minima model with a first minimum positioned at $\tilde{k}_{1}=1$ with $\hat{v}(\tilde{k}_{1})=-1$ and a second minimum at $\tilde{k}_{2}$ (see Table 2) with a varying vertical position, as it is shown in Fig. 5(a). The resulting phase diagram for this model is shown in Fig. 5(d). The obtained phase diagram presents the trivial homogeneous superfluid phase for low values of $\lambda^{2}U\rho$ and a phase transition from a hexagonal phase with characteristic wave vector $\tilde{k}_{1}$ (HS1) to one with characteristic wave vector $\tilde{k}_{2}$ (HS2) as $\hat{v}(\tilde{k}_{2})$ is decreased. Interestingly, the addition of a third resonant length scale of the $8$-QC phase to the pair potential $\hat{v}(k)$, either $k=\tilde{k}_{3}$ or $k=\tilde{k}_{4}$, in a scenario with two degenerate minima is not sufficient for the stabilization of the $8$-QC phase as occurs in the case of the decagonal pattern, see Fig. 4(e). Even in a scenario without any degeneracy we did not observe the stabilization of the $8$-QC pattern using only potentials with three characteristic length scales.
@@
In this way we proceed by allowing the existence of a fourth minimum in $\hat{v}(k)$. Hence, our pair potential potential have local minima at $k=\tilde{k}_{1}$, $\tilde{k}_{2}$, $\tilde{k}_{3}$ and $\tilde{k}_{4}$, see table 2. For this kind of model, we set $\hat{v}(\tilde{k}_{1})$ and $\hat{v}(\tilde{k}_{2})$ as degenerate minima, while we explore separately the phase diagram varying the vertical position of the other two minima present in $\hat{v}(k)$. Firstly, we set $\hat{v}(\tilde{k}_{4})=0$ and vary the depth of the minimum at $\tilde{k}_{3}$. The resulting model is presented in Fig. 5(b) and the corresponding phase diagram is shown in Fig. 5(e). Here we observe that in the low $\lambda^{2}U\rho$ regime, the system has a similar behavior to the two-minima model presented in Fig. 5(a), with the HS3 phase having characteristic momentum $\tilde{k}_{3}$. However, for large enough potential intensity we observe the stabilization of an $8$-QC in a narrow range of $\hat{v}(\tilde{k}_{3})$. Although, strictly speaking the $8$-QC exist as the ground state of this model, its stabilization occurs in a region where the validity of the mean field approximation itself can be questioned as the intensity of the pair potential in this region can be hardly considered as weak.
@@
In the last possible scenario, we explore a four-minima interaction potential fixing $\hat{v}(\tilde{k}_{3})=-0.6$, a value that in principle favors the formation of the octagonal quasicrystal without turning this minimum the dominant one, see Fig. 5(c). Meanwhile, we use the depth of the minimum at $k=\tilde{k}_{4}$ as running parameter for the construction of the phase diagram displayed in Fig. 5(f). As can be observed, this last family of models greatly enhance the stability of the $8$-QC phase. This is evident since, in this case, there is a significant range of $\hat{v}(\tilde{k}_{4})$ for which we have a direct transition from the homogeneous to the $8$-QC phase increasing the potential intensity. Moreover, the system naturally displays a HS4 phase with characteristic momentum close to $\tilde{k}_{4}$ when the potential minimum at this momentum is deep enough. The systematic study presented allow us to conclude that in order to stabilize the $8$-QC phase for the model considered, at least four characteristic wave vectors of this structure should be excited by the pair interaction potential $\hat{v}(k)$. Interestingly, the value of $\hat{v}(\tilde{k}_{4})$ seem to have a much stronger impact enhancing stability of the $8$-QC than the value of $\hat{v}(\tilde{k}_{3})$.
@@
Figure 6: (a) Real space mapping for the octagonal structure in the phase diagram of Fig. 5(f) for parameter values $\hat{v}(\tilde{k}_{4})=-1.25$ and $\lambda^{2}U\rho=0.50$ and its corresponding (b) diffraction pattern. The red circles indicates the two minima and the size of the Fourier modes in blue disks is scaled by a linear function, where the zero harmonic mode has been omitted. (c) Most relevant energy contributions to the total energy of the quasicrystalline structures as a contraction in the $\tilde{k}_{1}=1$ layer for the octagonal vectors. (d) Superfluid fraction for a fixed density value $\lambda^{2}U\rho=0.50$ in the octagonal quasicrystal for the ground state depicted in Fig. 5(f). The blue points represents the superfluid fraction obtained via mean field spectral variational method while the orange diamonds are determined via GPE simulations.
@@
A real space density configuration example and its corresponding Fourier transform plot, for the model in Fig. 5(c), are presented in Fig. 6(a) and (b), respectively. The parameters employed for the variational minimization are $\hat{v}(\tilde{k}_{4})=-1.25$ and $\lambda^{2}U\rho=0.50$. It is worth noticing how the stabilization of $8$-QC phase occurs when all the unstable momenta of the engineered pair potential are excited in the density pattern of the system. Moreover, the investigation of the energy behavior using a reduced Fourier expansion for the $8$-QC reveals a scenario similar to the one observed for the $10$-QC, where no exclusive form of contraction has a high impact on the energy values and, in this context, only the introduction of additional modes and the eventual emergence of new contractions with higher momentum are effective in the stabilization of the $8$-QC phase. We show in Fig. 6(c) an example of the leading contractions for this phase, the specific wave vectors employed for such contractions can switch depending on which is the most unstable wave vector in the system, however preserving the rule of having two modes excited at the most unstable momenta of the system, one at the second most unstable momenta and one with zero momentum.
@@
For completeness, we finally study the behavior of the superfluid fraction for the model described in Fig. 5(c) and Fig. 5(f), using $\hat{v}(\tilde{k}_{4})$ as running parameter and fixing $\lambda^{2}U\rho=0.50$. The results obtained are displayed in Fig. 6(d). As previously observed, since the HS1 phase does not excite Fourier modes with momentum $\tilde{k}_{4}$ or close to it, this phase is unaffected by changes in $\hat{v}(\tilde{k}_{4})$. In turn, when the system transition to a phase in which Fourier modes with momentum $\tilde{k}_{4}$, or close to it, are excited, to decrease the value of $\hat{v}(\tilde{k}_{4})$ promote localization of the density pattern, as explained in previous cases. This is the reason for the observed decreasing behavior of the superfluid fraction as we decrease the value of $\hat{v}(\tilde{k}_{4})$ for the 8-QC and the HS4 phase, respectively. Moreover, we can notice that as in previous cases, also the 8-QC phase is able to host significant superfluidity confirming the existence of a superfluid octagonal phase.

### v2

@@ III Dodecagonal quasicrystal
Moreover, we investigate how the superfluid properties of the $12$-QC phase are impacted by the shape and intensity of the pair interaction potential. We estimate the superfluid fraction by means of the Legget’s criterion and consider the scenario posed by Fig. 1(a). The superfluid fraction is calculated across a vertical line of the phase diagram of Fig. 1(d), specifically we consider $\gamma=0.50$ and use $\hat{v}(\tilde{k}_{2})$ as the running parameter. The results are shown in Fig. 2(d), where the blue points corresponds to the values of the superfluid fraction obtained from the density profile calculated using the mean field spectral variational method while the orange diamonds corresponds to the profile determined via the GPE solution. Within the HS1 phase, for $\hat{v}(\tilde{k}_{2})\gtrsim-1.5$, we observe a superfluid fraction essentially constant. This is explained by the fact that Fourier modes with $\lvert\bm{k}\rvert=\tilde{k}_{2}$ are not present in the HS1 state when the characteristic wave vector of the pattern is of the order of $\tilde{k}_{1}$. Moreover, the fact that superfluidity is significant along this phase allows us to conclude that it is actually a supersolid phase. On the other hand, in the region in which the model displays a $12$-QC phase, we can notice how the increase of the depth of the second minimum decreases the superfluid fraction. This occurs because decreasing the energy cost of the second minimum promotes higher Fourier amplitudes which implies in a more localized quasicrystalline density pattern, hindering superfluidity in the system. Nonetheless, despite the tendency of the superfluidity, within the quasicrystalline phase its value is still significant, allowing us to confirm that this phase is actually a superfluid dodecagonal phase. Finally, the expected decreasing behavior of the superfluid fraction as a function of the pair potential intensity $\gamma$, within each modulated phase, was also confirmed. The results for each quasicrystal structure can be found in Appendix B.
@@ IV Decagonal quasicrystal
Now we proceed to study the stabilization of the decagonal quasicrystal ($10$-QC) phase. We explore systematically pair potentials with multiple length scales in the form of several minima in $\hat{v}(k)$, positioned to enhance the stability of this structure. Naturally, the first model to be considered contains two minima, the first one at $k=\tilde{k}_{1}=1$ and the second one at $k=\tilde{k}_{2}$ (see Tab. 2). The corresponding pair interaction potential is displayed in Fig. 3(a), while the associated ground-state phase diagram is shown in Fig. 3(d). In this case, a phase transition is observed at moderate and high intensity of the pair potential $\gamma$ as the vertical position of the high momentum minima is varied. For low enough values of $\hat{v}(\tilde{k}_{2})$, the system displays a transition between two hexagonal solid phases in which the characteristic wave vector of the patterns changes from $\tilde{k}_{1}$ (HS1) to $\tilde{k}_{2}$ (HS2). Interestingly, despite the fact that this model produces the excitation of the two main wave vectors of the decagonal quasicrystal, this pair potential configuration is still favorable to the hexagonal structure even when the high momentum minimum dominates. This indicates the necessity of introducing additional minima in $\hat{v}(k)$ to achieve the stabilization of the decagonal quasicrystal pattern.
@@
To proceed we consider the case in which a third minimum in $k=\tilde{k}_{3}$ (see Tab. 2) is added to a two minima degenerate potential with $\hat{v}(\tilde{k}_{1})=\hat{v}(\tilde{k}_{2})=-1$. The resulting model for $\hat{v}(k)$ can be observed in Fig. 3(b) and the corresponding ground-state phase diagram using as running parameters the $\hat{v}(\tilde{k}_{3})$ and $\gamma$ is shown in Fig. 3(e). The alternative scenario in which the third minima is added in $\hat{v}(k)$ at $k=\tilde{k}_{4}$ is discussed in Appendix A. As can be noticed, the introduction of a third characteristic length scale allows the stabilization of a 10-QC phase for a certain range of $\hat{v}(\tilde{k}_{3})$. Interestingly, if $\hat{v}(\tilde{k}_{3})$ is deep enough the system reenters into a hexagonal phase (HS3) with characteristic wave vector $\tilde{k}_{3}$. Moreover, the quasicrystal phase exists up to the low intensity potential regime, making the system to present a direct transition from the homogeneous to the quasicrystal phase. However, this behavior is only present in a limited region of $\hat{v}(\tilde{k}_{3})$, raising naturally the question about what can be done to enhance further the stability or extension of this phase within the phase diagram.
@@
Figure 4: (a) Real space mapping for the decagonal structure in the phase diagram of Fig. 3(f) for parameter values $\hat{v}(\tilde{k}_{4})=-1.50$ and $\gamma=0.50$ and its corresponding (b) diffraction pattern. The red circles indicates the two minima and the size of the Fourier modes in blue disks is scaled by a linear function, where the zero harmonic mode has been omitted. (c) Most relevant energy contributions to the total energy of the quasicrystalline structures as a contraction in the $\tilde{k}_{1}=1$ layer for the decagonal vectors. (d) Superfluid fraction for a fixed intensity of the pair interaction, $\gamma=0.50$, for the the ground state phase diagram depicted in Figure 3(f). The blue points represents the superfluid fraction obtained via mean field spectral variational method while the orange diamonds are determined via GPE simulations.
@@
In order to answer this question, it is natural to consider the introduction of an additional length scale in the form of a minimum in $\hat{v}\left(k\right)$ at the remaining characteristic wave vector of the decagonal quasicrystalline structure. In this way, we add a minimum at $k=\tilde{k}_{4}$ (see Tab. 2 ) to a three minima potential that does not display a quasicrystal phase at any intensity of the pair potential. In terms of minima position, this model satisfy $\hat{v}(\tilde{k}_{1})=\hat{v}(\tilde{k}_{2})=-1$, $\hat{v}(\tilde{k}_{3})=0$ and $\hat{v}(\tilde{k}_{4})$ is taken as running parameter for the construction of the corresponding phase diagram. The typical form of this kind of pair potential is displayed in Fig. 3(c), while in Fig. 3(f) we present the corresponding ground-state phase diagram obtained. As can be observed, when the depth of the fourth minimum at $\tilde{k}_{4}$ is increased the system transitions from a HS1 to a $10$-QC phase. This means that only the addition of this extra minimum promotes the stabilization of the $10$-QC phase, but the extension of the quasicrystal phase is greatly enhanced. Further increasing the minimum depth leads to a transition from the decagonal quasicrystal to a hexagonal pattern with characteristic wave vector $\tilde{k}_{4}$, as expected when this minimum becomes the dominant one in the pair potential configuration.
@@
To deepen our understanding about the $10$-QC phase, we now focus on its structural properties. In Fig. 4(a) and (b), we present a typical density pattern configuration and its corresponding Fourier transform for a pair potential of the type presented in Fig. 3(c) ($\hat{v}(\tilde{k}_{4})=-1.50$ and $\gamma=0.50$). The radius of each circle in the Fourier transform plot is proportional to the corresponding Fourier amplitude. As we can see, the main modes excited in the density pattern corresponds to those promoted by the minima structure of the pair potential as in the case of the $12$-QC phase. An analysis of the potential energy contractions leading to dominant contributions, analogous to the one performed in the case of the dodecagonal structure, reveals that this structure does not have exclusive amplitude contractions which are specially impactful as in the previous case considered. In this case, contractions allowed for all quasicrystal structures, such as the one depicted in Fig. 4(c) for the decagonal pattern, are the leading contribution in the potential energy expansion.
@@
The behavior of the superfluid fraction for the model in Fig. 3(c), keeping $\gamma=0.50$ and using $\hat{v}(\tilde{k}_{4})$ as our running parameter, is shown in Fig. 4(d). We can observe that within the region corresponding to the HS1 phase, modifying $\hat{v}(\tilde{k}_{4})$ does not affect the superfluid fraction. As previously observed, since $\tilde{k}_{4}$ does not match any characteristic momentum of the hexagonal crystal solution with wave vector $\tilde{k}_{1}$, the hexagonal ground state is not affected by variations of $\hat{v}(\tilde{k}_{4})$ and hence neither the superfluid fraction. On the other hand, when the system transitions to the decagonal phase, we observe that making deeper the minimum at $\tilde{k}_{4}$ results in a decrease of the superfluid fraction. This behavior is a consequence of increasing the stability of the quasicrystalline phase, which results in an increase of the quasicrystal Fourier amplitudes. This in turn reflects on a more localized density profiles which hinder superfluidity. Moreover, is interesting to notice that close to the quasicrystal phase boundary, the superfluid fraction takes significant values, indicating the existence of a superfluid decagonal phase.
@@
Figure 5: (a)-(c) Pair interaction potential minima structure in Fourier space for the corresponding phase diagrams in (d)-(f), where the first length scale is positioned at $k=\tilde{k}_{1}=1$ and the secondary characteristic scales are presented in Tab. 2. Phase diagram for the minima structure associated with the octagonal quasicrystal where (d) $\bar{k}=\tilde{k}_{2}$ (e) $\bar{k}=\tilde{k}_{3}$ and (f) $\bar{k}=\tilde{k}_{4}$ indicates the position of the minimum in reciprocal space for which we modify the depth as a parameter to determine the ground state of the system. The notation employed to label the phases is analogous to the one established in Fig.1.
@@ V Octagonal quasicrystal
Finally, we focus on the study of the stability of the octagonal quasicrystal ($8$-QC) phase. As in the previous cases, we initially consider a two minima model with a first minimum positioned at $\tilde{k}_{1}=1$ with $\hat{v}(\tilde{k}_{1})=-1$ and a second minimum at $\tilde{k}_{2}$ (see Tab. 2) with a varying vertical position, as it is shown in Fig. 5(a). The resulting phase diagram for this model is shown in Fig. 5(d). The obtained phase diagram presents the trivial homogeneous superfluid phase for low values of $\gamma$ and a phase transition from a hexagonal phase with characteristic wave vector $\tilde{k}_{1}$ (HS1) to one with characteristic wave vector $\tilde{k}_{2}$ (HS2) as $\hat{v}(\tilde{k}_{2})$ is decreased. Interestingly, the addition of a third characteristic length scale of the $8$-QC phase to the pair potential $\hat{v}(k)$, either $k=\tilde{k}_{3}$ or $k=\tilde{k}_{4}$, in a scenario with two degenerate minima is not sufficient for the stabilization of the $8$-QC phase as occurs in the case of the decagonal pattern, see Fig. 4(e). Even in a scenario without any degeneracy we did not observe the stabilization of the $8$-QC pattern using only potentials with three characteristic length scales.
@@
In this way we proceed by allowing the existence of a fourth minimum in $\hat{v}(k)$. Hence, our pair potential potential have local minima at $k=\tilde{k}_{1}$, $\tilde{k}_{2}$, $\tilde{k}_{3}$ and $\tilde{k}_{4}$, see Tab. 2. For this kind of model, we set $\hat{v}(\tilde{k}_{1})$ and $\hat{v}(\tilde{k}_{2})$ as degenerate minima, while we explore separately the phase diagram varying the vertical position of the other two minima present in $\hat{v}(k)$. Firstly, we set $\hat{v}(\tilde{k}_{4})=0$ and vary the depth of the minimum at $\tilde{k}_{3}$. The resulting model is presented in Fig. 5(b) and the corresponding phase diagram is shown in Fig. 5(e). Here we observe that in the low $\gamma$ regime, the system has a similar behavior to the two-minima model presented in Fig. 5(a), with the HS3 phase having characteristic momentum $\tilde{k}_{3}$. However, for large enough potential intensity we observe the stabilization of an $8$-QC in a narrow range of $\hat{v}(\tilde{k}_{3})$. Although the $8$-QC phase is stabilized as the ground-state in this model, it only emerges in a regime of moderate to strong modulations. Taking into account that the implemented mean-field description considers a system in which particles are not localized, it is natural to seek a scenario where a direct transition from the homogeneous phase to the $8$-QC phase occurs. In such a case, one expects a significant degree of particle delocalization, thereby supporting the validity of the mean-field approximation. To this end, we now investigate an alternative configuration for the pair interaction potential, employing a different characteristic wave vector as the running parameter for our study.
@@
In this last scenario, we continue to explore a four-minima interaction potential, but now we fix $\hat{v}(\tilde{k}_{3})=-0.6$, a value that favors the formation of the octagonal quasicrystal without turning this minimum the dominant one, see Fig. 5(c). Moreover, we use the depth of the minimum at $k=\tilde{k}_{4}$ as a running parameter for the construction of the phase diagram displayed in Fig. 5(f). As can be observed, this last family of models greatly enhances the stability of the $8$-QC phase. This is evident since, in this case, there is a significant range of $\hat{v}(\tilde{k}_{4})$ for which we have a direct transition from the homogeneous to the $8$-QC phase increasing the potential intensity. Moreover, the system naturally displays a HS4 phase with characteristic momentum close to $\tilde{k}_{4}$ when the potential minimum at this momentum is deep enough. The systematic study presented allow us to conclude that in order to stabilize the $8$-QC phase for the model considered, at least four characteristic wave vectors of this structure should be excited by the pair interaction potential $\hat{v}(k)$. Interestingly, the value of $\hat{v}(\tilde{k}_{4})$ seem to have a much stronger impact enhancing stability of the $8$-QC than the value of $\hat{v}(\tilde{k}_{3})$.
@@
Figure 6: (a) Real space mapping for the octagonal structure in the phase diagram of Fig. 5(f) for parameter values $\hat{v}(\tilde{k}_{4})=-1.25$ and $\gamma=0.50$ and its corresponding (b) diffraction pattern. The red circles indicates the two minima and the size of the Fourier modes in blue disks is scaled by a linear function, where the zero harmonic mode has been omitted. (c) Most relevant energy contributions to the total energy of the quasicrystalline structures as a contraction in the $\tilde{k}_{1}=1$ layer for the octagonal vectors. (d) Superfluid fraction for a fixed value of the pair potential intensity, $\gamma=0.50$, for the ground state phase diagram in Fig. 5(f). The blue points represents the superfluid fraction obtained via mean field spectral variational method while the orange diamonds are determined via GPE simulations.
@@
A real space density configuration example and its corresponding Fourier transform plot, for the model in Fig. 5(c), are presented in Fig. 6(a) and (b), respectively. The parameters employed for the variational minimization are $\hat{v}(\tilde{k}_{4})=-1.25$ and $\gamma=0.50$. It is worth noticing how the stabilization of $8$-QC phase occurs when all the unstable momenta of the engineered pair potential are excited in the density pattern of the system. Moreover, the investigation of the energy behavior using a reduced Fourier expansion for the $8$-QC reveals a scenario similar to the one observed for the $10$-QC, where no exclusive form of contraction has a high impact on the energy values and, in this context, only the introduction of additional modes and the eventual emergence of new contractions with higher momentum are effective in the stabilization of the $8$-QC phase. We show in Fig. 6(c) an example of the leading contractions for this phase, the specific wave vectors employed for such contractions can switch depending on which is the most unstable wave vector in the system, however preserving the rule of having two modes excited at the most unstable momenta of the system, one at the second most unstable momenta and one with zero momentum.
@@
For completeness, we finally study the behavior of the superfluid fraction for the model described in Fig. 5(c) and Fig. 5(f), using $\hat{v}(\tilde{k}_{4})$ as running parameter and fixing $\gamma=0.50$. The results obtained are displayed in Fig. 6(d). As previously observed, since the HS1 phase does not excite Fourier modes with momentum $\tilde{k}_{4}$ or close to it, this phase is unaffected by changes in $\hat{v}(\tilde{k}_{4})$. In turn, when the system transition to a phase in which Fourier modes with momentum $\tilde{k}_{4}$, or close to it, are excited, to decrease the value of $\hat{v}(\tilde{k}_{4})$ promote localization of the density pattern, as explained in previous cases. This is the reason for the observed decreasing behavior of the superfluid fraction as we decrease the value of $\hat{v}(\tilde{k}_{4})$ for the 8-QC and the HS4 phase, respectively. Moreover, we can notice that as in previous cases, also the 8-QC phase is able to host significant superfluidity confirming the existence of a superfluid octagonal phase.